Hello,
Here is the question that I have to answer. I am totally lost!! I have read the book over and over again and cannot figure this out.
1. Give an example of an exponential function. Convert this exponential function to a logarithmic function. Plot the graph of both the functions.
I am guessing but I think that an example of an exponential function would be e^2-13=x. I have no idea how to then convert this to a logarithmic function and to plot a graph. Please someone help break this down for me to understand. I want to learn it!! Thank you in advance for your time.
Colleen

Answer by Alan3354(21611)

1. Give an example of an exponential function. Convert this exponential function to a logarithmic function. Plot the graph of both the functions.
I am guessing but I think that an example of an exponential function would be e^2-13=x. I have no idea how to then convert this to a logarithmic function and to plot a graph.
-------------------
x = e^2 - 13 is not a great example, since the right side is just constants.
y = e^x is better and simpler, or f(x) = e^x.
------------
If y = e^x, then ln(y) = x.
The base can be any constant, but e and 10 are most common, in fact "common logs" are base 10. e is used for some good reasons that you will be told when you are older (jk).
---------------------
To plot the graphs, dl the FREE software at
www.padowan.dk.com/graph/
If you email me via the thank you note, or at gsihoutx@aol.com, I can make the graphs and send them to you.